システム基礎講座 雑誌目録

平成 7 年 9 月 13 日

Artificial Intelligence  Vol. 65 (1994) – Vol. 71 No. 1+
Artificial Intelligence in Engineering  Vol.  9 (1994) – Vol.  9 No. 1+
Artificial Intelligence Review  Vol.  8 (1994) – Vol.  8 No. 1+
Complex Systems  Vol.  8 (1994) – Vol.  8 No. 4+
Computational Intelligence  Vol. 10 (1994) – Vol. 10 No. 4+
Computational Optimization and Applications  Vol.  2 (1993) – Vol.  3 No. 4+
IEEE Transactions on Neural Networks  Vol.  4 (1993) – Vol.  5 No. 6+
Immunological Reviews  No.  143 (1995) +
Mathematical Programming  Vol. 58 (1993) – Vol. 67 No. 3+
Mathematics of Operations Research  Vol. 18 (1993) – Vol. 19 No. 3+
Neural Networks Vol. 6 (1993) – Vol. 7 No. 5+
Operations Research  Vol. 41 (1993) – Vol. 42 No. 5+
Operations Research Letters  Vol. 13 (1993) – Vol. 16 No. 4+
ORSA Journal on Computing Vol. 5 (1993) – Vol. 6 No. 2+
SIAM Journal on Control and Optimization  Vol. 31 (1993) – Vol. 32 No. 6+
SIAM Journal on Numerical Analysis  Vol. 30 (1993) – Vol. 31 No. 5+
SIAM Journal on Optimization Vol. 3 (1993) – Vol. 4 No. 3+
SIAM Journal on Scientific Computing  Vol. 14 (1993) – Vol. 15 No. 5+
The Mathematica Journal Vol. 5 (1995) – Vol. 5 No. 1+
Transportation Research. Part B: Methodological Vol. 27B (1993) – Vol. 28B No. 4+
Transportation Science  Vol. 27 (1993) – Vol. 28 No. 2+
応用数理 Vol. 3 (1993) – Vol. 4 No. 3+
コンピュートロール  No. 12,26,36,37
数理科学  Vol. 32 (1994) – Vol. 33 No. 1+
bit  Vol. 26 (1994) – Vol. 26 No. 12+
bit 別冊  (1994) 1 – (1994) 8

[Annals of Operations Research]

Vol. 18 Facility Location Analysis: Theory and Applications,

F.V. Louveaux, M. Labbé, J.F. Thisse (eds.), 1989. Vol. 25 Computational Methods in Global Optimization,

P.M. Pardalos, J.B. Rosen (eds.), 1990. Vol. 30 Stochastic Programming – Part I,

J.R. Birge, R.J.-B. Wets (eds.), 1991. Vol. 31 Stochastic Programming – Part II,

J.R. Birge, R.J.-B. Wets (eds.), 1991. Vol. 33 Topological Network Design,

J. MacGregor Smith, P. Winter (eds.), 1991. Vol. 34 Hierarchical Optimization,

G. Anandalingam, T.L. Friesz (eds.), 1992. Vol. 40 Locational Decisions,

Z. Drezner (ed.), 1992. Vol. 41 Tabu Search,

F. Glover, M. Laguna, E. Taillard, D. de Werra (eds.), 1993. Vol. 43 Applied Mathematical Programming and Modelling,

G. Mitra, I, Maros (eds.), 1993. Vol. 44 Advances in Equilibrium Modeling, Analysis and Computation,

A. Nagurney (ed.), 1993. Vol. 45 Financial Engineering,

H. Konno, D.G. Luenberger, J.M. Mulvey (eds.), 1993.

Vols. 46/47 Degeneracy in Optimization Problems, T. Gal (ed.), 1993.